### v2

# Bioverse: GMT and ELT Direct Imaging and High-Resolution Spectroscopy Assessment – Surveying Exo-Earth O2 and Testing the Habitable Zone Oxygen Hypothesis

Abstract
Biosignature detection in the atmospheres of Earth-like exoplanets is one of the most significant and ambitious goals for astronomy, astrobiology, and humanity. Molecular oxygen is among the strongest indicators of life on Earth, but it will be extremely difficult to detect via transmission spectroscopy. We used the Bioverse statistical framework to assess the ability to probe Earth-like O2 levels on hypothetical nearby habitable zone exo-Earth candidates (EECs) using direct imaging and high-resolution spectroscopy on the Giant Magellan Telescope (GMT) and the Extremely Large Telescope (ELT). Assuming continued improvement in instruments and data processing, our analysis highlights the best-case scenarios. Earth-like O2 levels could be probed on up to $\sim$7 and $\sim$19 EECs orbiting bright M dwarfs within 20 pc in a hypothetical 10-year survey on the GMT and ELT, respectively. Four known super-Earth candidates, including Proxima Centauri b, could be probed for O2 within about one week of observations on the ELT and a few months on the GMT. We also assessed the ability of the ELT to test the habitable zone oxygen hypothesis – that habitable zone Earth-sized planets are more likely to have O2 – within a 10-year survey using Bioverse. Testing this hypothesis requires either $\sim$1/2 of the EECs to have O2 or $\sim$1/3 if $\eta_{\oplus}$ is large. A northern hemisphere large-aperture telescope, such as the Thirty Meter Telescope (TMT), would expand the target star pool by about 25%, reduce the time to probe biosignatures on individual targets, and provide an additional independent check on potential biosignature detections.
Exoplanet systems (484) — Exoplanets (498) — Exoplanet Atmospheres (487) — Biosignatures (2018) — Astrobiology (74) — Habitable Zone (696) — Habitable Planets (695) — Astronomical Simulations (1857) — Bayesian Statistics (1900) — Parametric Hypothesis Tests (1904)
software: astroplan (Morris et al., 2018), astropy (Astropy Collaboration et al., 2013, 2018, 2022), Bioverse (Bixel & Apai, 2021; Hardegree-Ullman et al., 2023; Schlecker et al., 2024), dynesty (Speagle, 2020), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), pandas (Wes McKinney, 2010; pandas Development Team, 2020), PyAstronomy (Czesla et al., 2019)
1 Introduction
Methods for probing the atmospheres of exoplanets have matured to the point that we are within reach of biosignature detection on rocky worlds. Transmission spectroscopy has become a powerful tool to characterize the atmospheres of planets from the ground (e.g., Snellen et al., 2008, 2010; Stevenson et al., 2014), and from space with HST (e.g., Charbonneau et al., 2002; Sing et al., 2016; Zhang et al., 2018) and now JWST (e.g., JWST Transiting Exoplanet Community Early Release Science Team et al., 2023; Moran et al., 2023). In the search for biosignatures via transmission spectroscopy, JWST might prove fruitful in the detection of species such as CH4, H2O, and CO2 (e.g., Wunderlich et al., 2019), however, it will not be able to detect O2 (Lustig-Yaeger et al., 2019; Wunderlich et al., 2019; Pidhorodetska et al., 2020), a major byproduct of the biological process of oxygenic photosynthesis on Earth. Snellen et al. (2013), Rodler & López-Morales (2014), Serindag & Snellen (2019), López-Morales et al. (2019), Currie et al. (2023), and Hardegree-Ullman et al. (2023) studied the prospects of detecting O2 via transmission spectroscopy using upcoming 25–40-meter class telescopes such as the Giant Magellan Telescope (GMT, Johns et al., 2012), Thirty Meter Telescope (TMT, Nelson & Sanders, 2008; Sanders, 2013), and European Southern Observatory Extremely Large Telescope (ELT, Gilmozzi & Spyromilio, 2007). Hardegree-Ullman et al. (2023) took into account practical ground-based observing constraints such as transit observability and relative system velocities to mitigate telluric blending. They concluded that even in a very optimistic observing scenario where signals from the GMT, TMT, and ELT and all observable transits could be combined, it will likely take decades to make an O2 detection on a single transiting habitable zone Earth-sized exoplanet.
[…]
1.1 Previous Direct Imaging Studies
There have been numerous studies investigating the potential for a next-generation space telescope (e.g., LUVOIR, HabEx, LIFE, The LUVOIR Team, 2019; Gaudi et al., 2020; Quanz et al., 2022) to probe biosignatures in reflected light with direct imaging and low-resolution spectroscopy (e.g., Feng et al., 2018; Bixel & Apai, 2021; Damiano & Hu, 2022; Konrad et al., 2022; Robinson & Salvador, 2023; Susemiehl et al., 2023). Due to the recommendations brought forth in the Astro2020 Decadal Survey (NASEM, 2021), the Habitable Worlds Observatory (HWO) is being developed with the goal to identify and directly image at least 25 potentially habitable worlds using a telescope about the same size as JWST and optimistically launching in the 2040s. In the meantime, the GMT, TMT, and ELT are expected to be operational in the early 2030s. The significantly larger apertures of these telescopes will be able to probe smaller inner working angles and study the habitable zones of nearby M dwarfs, which will be inaccessible to HWO.
Ground-based spectroscopy of biosignatures necessitates the use of high resolution spectrographs ($R\gtrsim 50,000$) to allow separation of Earth’s telluric lines from those of the exoplanet atmosphere, assuming sufficient line Doppler shifts to minimize line blending (see, e.g., Rodler & López-Morales, 2014; López-Morales et al., 2019, for examples of different resolutions and their effect on line blending). Sparks & Ford (2002) proposed combining a coronographic imager to a high spectral resolution integral-field spectrograph for detection and characterization of exoplanets. As Snellen et al. (2015) showed, the achievable contrast of a combined high-contrast imaging and high-resolution spectrograph system is the product of the individual system achievable contrasts. For example, if each individual system can reach a contrast of $\sim$10-5, the combined system could reach contrasts of $\sim$10-10 (see additional discussion in Section 2.2).
[…]
Snellen et al. (2015) investigated using an optical high-resolution spectrograph (R=100,000) in combination with an extreme adaptive optics (AO) system on the ELT to observe a hypothetical 1.5 $R_{\oplus}$ planet in the habitable zone of Proxima Centauri in reflected light (prior to the discovery of Proxima Centauri b by Anglada-Escudé et al., 2016). They concluded that such a planet would be detectable in broadband (0.6–0.9 $\mu$m) reflected light in 10 hours of integration time at a S/N of $\sim$10 after cross-correlating the hypothetical observed spectrum with a model template spectrum, but they did not explore specific biosignature detection.
Since its discovery, most of the focus on high-contrast imaging and high-resolution reflected-light spectroscopy has been on Proxima Centauri b. This is likely driven by the planet’s location within the habitable zone, and a measured minimum mass ($M\sin i$) likely between $1.0~{}M_{\oplus}$ (Damasso et al., 2020) and $1.27~{}M_{\oplus}$ (Anglada-Escudé et al., 2016). Bixel & Apai (2017) placed probabilistic mass and radius constraints on Proxima Centauri b at $M=1.63^{+1.66}_{-0.72}~{}M_{\oplus}$ and $R=1.07^{+0.38}_{-0.31}~{}R_{\oplus}$, which is consistent with a rocky composition, but there is still a $\sim$10chance the planet has a significant amount of ice or a volatile envelope. We do not yet have enough information to truly constrain the nature of Proxima Centauri b, so assuming it is Earth-like or an Earth-analog is an optimistic assumption.
[…]
Wang et al. (2017) simulated observations in reflected light with low to high-resolution spectra of Proxima Centauri b (assuming it is an Earth-like planet) and a hypothetical Earth-like habitable zone planet orbiting an M dwarf at 5 pc using a 30-meter class telescope at near-infrared bands. They used near-infrared (1–2.5 $\mu$m) R=1,000–100,000 Earth-like exoplanet spectra, considering atmospheric chemistry and using a radiative transfer model from Hu et al. (2012a, b, 2013); Hu & Seager (2014). For a 100-hour integration time, their simulations yielded S/N$>$10 for H2O, O2, CO2, and CH4, typically when R$>$10,000 and star-light suppression was $>10^{-7}$. These simulations simplistically addressed line blending, assuming the radial velocity of the exoplanet shifts its atmospheric signature by tens of km s-1. An updated version of this analysis was provided by Zhang et al. (2024), where they simulated direct imaging of 10 nearby rocky planets at R=1000 with HARMONI-like and METIS-like instruments on the ELT in search of biosignatures. They indicated that CO2, CH4, and H2O should be detectable on GJ 887 b and Proxima Centauri b at a S/N$>$5 with integration times on the order of $<$100 hours.
Hawker & Parry (2019) focused on the detectability of Earth-like levels of O2 on Proxima Centauri b with high-resolution spectroscopy and high contrast imaging on the ELT using HIRES and HARMONI-like instruments. They concluded a S/N=3–5 detection could be obtained in 30–70 hours of integration time. More recently, Vaughan et al. (2024) simulated observations of Proxima Centauri b using HARMONI on the ELT at its highest resolving power of R=17,385. Their simulations indicated an atmosphere on Proxima Centauri b (assuming it is Earth-like) could be characterized at S/N=5 in 20–30 hours of integration time. These observations would be particularly sensitive to CH4, but CO2 would be harder to detect.
While these previous studies have made significant contributions to assessing direct imaging capabilities for individual cases, they were intrinsically constrained in their scope for broader hypothesis testing. Our study seeks to build upon the above foundational efforts by employing the Bioverse framework, which allows us to go beyond previous studies by leveraging exoplanet demographics information, and performing realistic survey simulations and hypothesis testing.
1.2 Bioverse
Bioverse was developed by Bixel & Apai (2021) as a modular framework to generate and classify exoplanets based on exoplanet demographics, simulate exoplanet surveys for upcoming telescopes and missions, and compute the diagnostic power of future surveys in testing population-level hypotheses. Bioverse folds in state-of-the-art exoplanet occurrence rate calculations, incorporates a complete nearby star catalog out to $\sim$100 pc (Gaia red-band magnitude range $19.90.5<G_{RP}<$ 19.9), and accounts for realistic ground-based observing constraints, allowing exploration of the full potential of a broad statistical survey of nearby exoplanets.
The general workflow of Bioverse starts with generating synthetic populations of stars with planetary systems and to injecting a statistical trend into that population. Then, a simulated survey “observes” a sample of the synthetic planets and collects measurement data according to projected uncertainties and survey strategies. Next, the hypothesis testing module allows quantification of how well the injected trend is recovered, and how strongly parameters of the model shaping the trend can be constrained. This procedure is typically repeated in a Monte Carlo fashion, varying astrophysical unknowns or features of the survey, which allows for trade studies on different survey designs.
After introducing the framework in Bixel & Apai (2021), Hardegree-Ullman et al. (2023) added a stellar catalog based on Gaia Data Release 3, the capability to determine transit observability for ground-based large-aperture telescopes, and simulations for oxygen detection via transmission spectroscopy. Schlecker et al. (2024) introduced selectable mass-radius relations, a mission simulator for ESA’s PLATO mission (Rauer et al., 2016), and a runaway greenhouse climate model. Here, we built upon these previous studies to assess the direct imaging and high resolution spectroscopy potential for reflected light studies with the GMT and ELT, specifically focusing on the ability to probe Earth-like O2 levels. Leveraging all modules of Bioverse, we seek to test the “habitable zone oxygen hypothesis,” i.e., the hypothesis that Earth-sized planets within a star’s habitable zone are more likely to have Earth-like O2 levels. This will inform us if a hypothetical 10-year survey of EECs with the ELT will be sufficient to test this hypothesis.
In Section 2 we outline the direct imaging observing considerations and instrument models. We simulate a 10-year survey with the GMT and ELT to probe Earth-like levels of O2 on habitable zone Earths and super-Earths in Section 3, and follow with testing the habitable zone oxygen hypothesis in Section 4. In Section 5 we discuss the implications of our simulations, and we summarize and conclude our results in Section 6.
2 Observing Considerations
Earth has evolved from a methane and carbon dioxide-rich atmosphere over 2.4 Gyr ago to its present-day oxygen-rich state due to oxygenic photosynthesis (e.g., Lyons et al., 2021). While it is likely that habitable zone Earth-sized exoplanets have atmospheres with diverse compositions and in different states of evolution, we adopt present-day Earth as a starting point to simplify simulations. The following simulations are focused on probing present-day Earth-like atmospheric conditions on rocky, habitable-zone exoplanets.
2.1 Planet Generation and Observational Constraints
In order to provide a realistic survey simulation, we used Bioverse to generate a set of exoplanets orbiting bright stars (Gaia red-band $G_{RP}\leq 10$) out to 20 pc from the Sun based on planet occurrence rates from Bergsten et al. (2022), which was first introduced to Bioverse in Schlecker et al. (2024). The underlying parametric model in Bergsten et al. (2022) is derived from planet radius, orbital period, and stellar host mass distributions in Kepler’s FGK and early M stars sample, considering planets with radii up to $3.5~{}R_{\oplus}$ and orbiting out to 100 days. The original SAG 13-based (Kopparapu et al., 2018) planet occurrence rate module introduced by Bixel & Apai (2021) was derived from a compilation of Kepler planet occurrence rates prior to 2017, and did not include a stellar mass dependence. Both of the planet occurrence rate modules are model-dependent for planets orbiting beyond 100 days. This does not affect our study much because an overwhelming majority of targets in our simulated surveys are M dwarfs (see Figure 5) with habitable zones closer than 100 days. We note that the SAG 13 occurrence rates are more optimistic than the Bergsten et al. (2022) values by a factor of $\sim$2, but we opt to use the Bergsten et al. (2022) values in this study because it is not only more recent, but also from a uniform study rather than a compilation.
From the set of generated exoplanets, we only considered planets that meet the criteria of an exo-Earth candidate (EEC): the planet is within the conservative habitable zone boundaries between runaway and maximum greenhouse as defined by Kopparapu et al. (2014), and the planet radius falls between the range $0.8S^{0.25}<R_{p}<1.4~{}R_{\oplus}$, where $S$ is stellar incident flux in units of present-day Earth insolation flux. The lower limit is the theoretical minimum planet size needed to retain an atmosphere (Zahnle & Catling, 2017), and was previously adopted for use in Bioverse by Bixel & Apai (2021) and Hardegree-Ullman et al. (2023).
In order to generate realistic surveys, for each EEC we simulated one Earth year of observations at one hour intervals and computed practical observing constraints based on the locations of the GMT and ELT. Simple constraints such as target observability at night and airmass between $z=1.0$ and 2.0 were modeled using astroplan (Morris et al., 2018). When a planet was observable, we computed relative radial velocities and planet-to-star contrast ratios.
Since we have to observe O2 on exoplanets through the O2-rich atmosphere of Earth, we must account for relative radial velocities ($\Delta$RV) in order to mitigate line blending effects. Relative radial velocities are given by:
$\Delta\mathrm{RV}=\mathrm{RV}_{\star}+\mathrm{RV}_{\mathrm{orb}}-\mathrm{RV}_{\mathrm{bary}},$ (1)
where $\mathrm{RV}_{\star}$ is the radial velocity of the host star, $\mathrm{RV}_{\mathrm{orb}}$ is the orbital velocity of the exoplanet with respect to its host star, and $\mathrm{RV}_{\mathrm{bary}}$ is the velocity of the observer on Earth with respect to the solar system barycenter. The host star radial velocity comes from the Bioverse stellar catalog (Hardegree-Ullman et al., 2023), which is based on Gaia Data Release 3. The uncertainties of these radial velocities are on the order of a few km s-1 (Gaia Collaboration et al., 2021).
Orbital velocity was computed from:
$\mathrm{RV}_{\mathrm{orb}}=\sqrt{GM_{\star}\left(\frac{2}{r}-\frac{1}{a}\right)},$ (2)
where $G$ is the gravitational constant, $M_{\star}$ is the mass of the host star, $r$ is the distance of the exoplanet to the host star at the time of observation (computed in Bioverse using Newton’s method to numerically solve Kepler’s equation), and $a$ is the semi-major axis of the exoplanet. Observer barycentric velocity at the time of observation was computed using the baryCorr function from the PyAstronomy package (Czesla et al., 2019). For an $R=100,000$ spectrograph observing the O2 A-band, severe line blending occurs when $|\Delta\mathrm{RV}|<13$ km s-1 and $30\ \mathrm{km\ s}^{-1}<|\Delta\mathrm{RV}|<55\ \mathrm{km\ s}^{-1}$ (López-Morales et al., 2019; Hardegree-Ullman et al., 2023).
For direct imaging observations, the projected distance of the exoplanet from its host star ($\theta$ [arcsec] = $r$ [au]/$d$ [pc], where $d$ is the distance to the system) must be larger than the inner working angle of the observing instrument. The diffraction limit of a telescope is $1.22\lambda/D$. In practice, we set the observable inner working angle to be greater than $2\lambda/D$ (e.g., Lovis et al., 2017; Walter et al., 2020). At the wavelength of the O2 A-band (760 nm), $2\lambda/D$ is 13 mas, 11 mas, and 9 mas for the GMT, TMT, and ELT, respectively.
In order to observe the exoplanet, the contrast ratio between the planet and the star must be above the detection limit of the instrument. This contrast ratio is given by:
$C=\frac{F_{p}(\lambda,\alpha)}{F_{\star}(\lambda)}=A_{g}(\lambda)g(\alpha)\left(\frac{R_{p}}{a}\right)^{2},$ (3)
where $F_{p}(\lambda,\alpha)$ is the exoplanet flux at wavelength $\lambda$ and phase angle $\alpha$, $F_{\star}(\lambda)$ is the host star flux, $A_{g}$ is the geometric albedo, $g(\alpha)$ is the scattering phase function, $R_{p}$ is the exoplanet radius, and $a$ is the exoplanet semi-major axis. Our models assume a Lambertian phase function:
$g(\alpha)=\frac{\sin\alpha+(\pi-\alpha)\cos\alpha}{\pi},$ (4)
[…]
2.2 Instrument Models
In order to provide a realistic simulation of extreme-AO performance, we need models for planned instruments, which currently include GMagAO-X (Males et al., 2022) for the GMT and the Planetary Camera and Spectrograph (PCS, Kasper et al., 2021) for the ELT. However, full end-to-end simulations of these high-order AO systems are very expensive in terms of computational time and complexity. Therefore, we used a semi-analytical approach to the performance simulation (Jolissaint, 2010). The semi-analytical approach assumes that all temporal evolution of the turbulence is driven by Taylor’s frozen flow. The frozen flow approximates the short time evolution by shifting the spatial structure of the turbulence with single-wind velocity. The time evolution of any parameter, $f$ for example, can then be replaced by a shifted version, $f(\vec{r},t)=f(\vec{r}-\vec{v}t)$. This time evolution underpins the semi-analytical model and allows us to express an adaptive optics system as a series of spatial filters that act on the turbulence Power Spectral Density (PSD). The spatial frequency approach has been verified extensively in comparison with full end-to-end models (Males et al., 2021). The AO system is controlled using integral control, $\mathrm{DM}_{i+1}=\mathrm{DM}_{i}-g\varepsilon_{i}$. Here, $\varepsilon_{i}$ is the wavefront error at time step $i$ and $\mathrm{DM}_{i}$ is the DM shape at time step $i$. The feedback gain $g$ is a function of spatial frequency and was optimized following Gendron & Léna (1994); Jolissaint (2010).
We simulated two different representative optimizations of the AO system. In the first, the AO system was allowed to run as fast as possible, which is 3 kHz for both GMagAO-X and PCS. The gain for each spatial frequency was optimized at this AO loop speed. In the second approach, we optimized the Strehl by changing both the AO loop speed and modal gains. These two approaches were used because optimizing for only Strehl does not necessarily lead to the best contrast. The contrast at small angular separations is driven by the servo lag error. This error term is a larger contributor on GMT/ELT-sized instruments than on current instruments (Males & Guyon, 2018). Therefore, it is possible to increase the contrast at smaller angular separations by running the AO system faster at the cost of lower Strehl. This leads to a better final signal-to-noise ratio. The two different approaches are called ‘Max Speed’ and ‘Strehl Optimization’ in this work. A more thorough assessment of which instrument and observing modes were selected for scientific analysis in Section 3 is given in Appendix A.
AO performance strongly depends on the atmospheric conditions assumed. For our simulations, we followed Prieto et al. (2010) and Thomas-Osip et al. (2010) for the GMT (see Table 1). The first quartile conditions have a Fried parameter of $r_{0}=0.22$ m and a $C_{n}^{2}$ integrated wind velocity of 9.4 m s-1. Here $C_{n}^{2}$ is the refractive index structure parameter. The second quartile conditions are $r_{0}=0.17$ m and a mean wind speed of 18.4 m s-1. For the ELT, we followed Sarazin et al. (2013): the first and second quartile conditions are $r_{0}=0.24$ m with a velocity of $10$ m s-1 and $r_{0}=0.15$ m with a velocity of 11 m s-1, respectively.
The design of GMagAO-X is currently further along than that of the other extreme-AO systems (Males et al., 2024; Haffert et al., 2024). Therefore, we used the current design parameters for GMagAO-X as reference. Specifically, we assumed 188 actuators across the pupil and a pyramid wavefront sensor. The design of PCS is currently still in its conceptual phase and no hard decisions have been made (Kasper et al., 2021). Therefore, we considered two different DM sizes for PCS: 200 actuators across the pupil based on the proposed density for EPICS (Kasper et al., 2010), and 128 actuators across the pupil based on a more recent ELT extreme-AO design (Madec et al., 2022).
The output of the semi-analytical model is a post-AO PSD. The PSD still has to be propagated through an actual coronagraphic system to measure the amount of residual scattered star light. The expected exposure times of our simulations are on the order of hours. This is much longer than the typical speckle lifetime of 1 to 10 ms (Males et al., 2021). We can treat the different spatial frequencies as incoherent because of the many orders of magnitude differences in characteristic timescales. With that in mind, the propagation through a coronagraph can be viewed as an incoherent sum of plane waves (Sauvage et al., 2010; Herscovici-Schiller et al., 2017). The theoretical perfect coronagraph removes only the piston mode from the incoming wavefront (Guyon et al., 2006; Deshler et al., 2024). However, such coronagraphs are extremely sensitive to residual jitter and partially-resolved stars (Guyon et al., 2006). Therefore, we used a 4th-order perfect coronagraph that also removes the tip/tilt modes from the wavefronts beside the piston mode. This increases the inner-working angle of the coronagraph and it increases the system robustness against stellar angular diameter and residual tip/tilt jitter(Guyon et al., 2006; Belikov et al., 2021). Given the size of the GMT/ELT, it will not be realistic to work with a 2nd-order coronagraph that only removes pistons. Finally, the effects of time varying quasi-static speckles are ignored in the simulations for two reasons. For high-resolution spectroscopy, the only thing that matters is the amount of starlight that is left in the final spectra. The cross-correlation function is differentiating between stellar residuals and planet light through spectral information only and the associated noise is photon noise of the time-integrated spectrum intensity (Snellen et al., 2015). The speckle lifetime affects the variance of this intensity but not the mean (shorter lifetime $\rightarrow$ smaller variance/mean). We assume that the quasi-static speckles can be controlled and removed to a level that is below the halo of residual atmospheric speckles using newly implemented focal plane wavefront control techniques (Potier et al., 2022; Haffert et al., 2023; Galicher et al., 2024) and that means the photon noise from the atmospheric speckles will dominate the noise budget. Therefore, the total power of the low-spatial frequency content of the quasi-static speckles was set to 1 nm and the high-spatial frequency content was set to 25 nm and added to our simulations. This amount is similar to what has been achieved with MagAO-X (Van Gorkom et al., 2021). We also assume that the efficiency of quasi-static speckle removal process will continue to improve over the next few years. An example output of the semi-analytical modeling is shown in Figure 1.
Figure 1: Output of the semi-analytical AO simulations for ELT/PCS with the 128 actuators model. The left figure shows the simulated PSF and the right figure shows the coronagraphic stellar image. There are two visible circular regions. The largest corresponds to the control radius of the AO system and the smaller one corresponds to the control radius of the non-common path aberrations. The contrast within the inner radius is dominated by the wind-driven halo. Axes are shown in both mas and $\lambda/D$, assuming $\lambda=765$ nm and $D=39.3$ m.
Figure 2: Coronagraphic encircled energy throughput ($1.8\lambda/D$ diameter) as a function of angular separation for an 8th magnitude star. A 4th-order perfect coronagraph with the ELT aperture was used in these simulations. The top row shows the PSFs corresponding to the four orange points on the throughput curve from 0 mas (left) to 20 mas (right) separation. The PSF is deformed close to the inner-working angle of the coronagraph.
ELT Parameters
Telescope area 978 m2
Number of actuators 128 / 200
Number of wfs pixels 125,000
GMT Parameters
Telescope area 358 m2
Number of actuators 188
Number of wfs pixels 111,000
Shared AO System Parameters
Zeropoint flux in I band 4520 photons s-1 cm-2 nm-1
Bandwidth I band 150 nm
WFS wavelength 765 nm
Loop speed 250 - 4000 Hz
Intrinsic delay 1.5 frames
Computational delay 250 $\mu$s
Throughput to WFS camera 10 %
WFS detector variance 2.0 $e^{-}$
Photon noise sensitivity $1/\sqrt{2}$
Read noise sensitivity $1/\sqrt{2}$
Common path throughput to spectrograph 25 %
Coronagraph 4th order perfect coronagraph (Guyon et al., 2006)
Site Parameters
$r_{0}$ 10 - 20 cm
$L_{0}$ 25 m
$v$ 15 - 40 m s-1
Spectrograph Parameters
read noise $1e^{-}$ pixel-1
dark current $10^{-4}$ $e^{-}$ s-1 pixel-1
Spectrograph throughput 35 %
Detector quantum efficiency 86 %
Science wavelength 765 nm
Spectral bandwidth 10 nm
Resolving power 100,000 – 500,000
Spectral resolution 3 channels per resolving element
Pixels per channel 2 pixels
Table 1: Parameters used to calculate S/N (Equation 5).
In order to calculate the signal-to-noise ratio of the high-contrast, high resolution spectroscopy observations of the exoplanet atmosphere O2 A-band, we follow the basic principle of Snellen et al. (2015) and Lovis et al. (2017):
{widetext}
$S/N=\frac{\eta\Delta tT_{p}T_{I}T_{AO}T_{QE}CF_{\star}}{\sqrt{\Delta tT_{\star}T_{I}T_{AO}T_{QE}F_{\star}+\Delta t\sigma^{2}_{\mathrm{sky}}+N_{\mathrm{exp}}N_{\mathrm{pixels}}\sigma^{2}_{\mathrm{read}}+N_{\mathrm{pixels}}\Delta t\sigma^{2}_{\mathrm{dark}}}},$ (5)
[…]
At high resolution ($R\geq 100,000$), we can mitigate Earth’s telluric signal by avoiding observing the exoplanet at relative radial velocities at which line blending is severe (Section 2.1). Disentangling the star and exoplanet signal happens in post-processing by cross-correlating the reflected-light spectrum with the stellar spectrum. Cross-correlation post-processing efficiency ($\eta$) for the O2 A-band converges to 95% for $R\geq 100,000$ (Fowler et al., 2023) if the planet is sufficiently Doppler-shifted (Hardegree-Ullman et al., 2023). Rearranging Equation 5, we computed the achievable contrast $C$ at S/N=5 for 1-hour total integration times (assuming four 15-minute exposures) for different host star apparent magnitudes and angular separations for the GMT (Table 3) and ELT (Table 4). The brightest stars might require shorter exposures, but the dominant detector noise will be dark current, which scales with exposure time, rather than read noise, so the effect is insignificant.
Figure 3: The S/N after a 1 hour integration for different planet contrasts and stellar magnitudes. The S/N=5 detection limit is shown by the dashed grey line.
[…]
2.3 Time Needed to Probe O2
With our new instrument model, we now have to determine how long it will take to probe O2. For each simulated exoplanet, we first computed the total number of hours the planet would be observable from the GMT and ELT locations for one year of observations ($N_{\text{hr-obs}}$) accounting for the aforementioned ground-based observing limitations and exoplanet orbital properties. Since the modeled planets change illumination and scattering phase throughout their orbits, we computed the contrast ratio from Equation 3 at each observable hour and used the resulting median contrast ratio ($\widetilde{F_{p}/F_{\star}}$) for our calculations. Tables 3 and 4 show the expected achievable post-processing contrasts from the instrument models in Section 2.2 for the GMT and ELT, respectively, at S/N = 5 for a one hour integration of the O2 A-band at $R=100,000$ for different host magnitudes and angular separations.
It would be ideal to optimize observations based on maximizing Equation 3. However, this depends on several factors, including well constrained exoplanet orbits and phases and appropriate relative system velocities coinciding with Q1 or Q2 observing windows. It is impractical to assume observations will only happen when all these factors are most optimal, so we opted for a more conservative median observing situation. For each simulated exoplanet, we interpolated the host star $G_{RP}$ magnitude and median angular separation ($\tilde{\theta}$) to these tables to yield a 1-hour contrast ($C_{1\text{hr}}$). Total integration time to achieve a S/N = 5 detection of Earth-like levels of O2 on the simulated planet ($t_{\text{hr-}5\sigma}$) is given by:
$t_{\text{hr-}5\sigma}=\left(\frac{C_{1\text{hr}}}{\widetilde{F_{p}/F_{\star}}}\right)^{2}.$ (6)
To translate total integration time to years of observations, we need to account for both target observability ($N_{\text{hr-obs}}$) and acceptable site observing conditions ($Q$):
$t_{\text{yr-}5\sigma}=Q\frac{t_{\text{hr-}5\sigma}}{N_{\text{hr-obs}}},$ (7)
where $Q=4$ for first quartile observing conditions (best observing site weather, ideal for extreme-AO, $\sim$25of telescope observing time, will take four times longer to achieve S/N) and $Q=2$ for second quartile observing conditions (best to average observing site weather, $\sim$50of telescope observing time, will take twice as long to achieve S/N). It is worth emphasizing that $N_{\text{hr-obs}}$ only counts times at which the relative system velocities yield unblended lines and the planet is observable from the observing site. We simulate first and second quartile observing conditions for each instrument mode, and assess the trade-offs of different observing scenarios in the Appendix.
GMT, Max Speed, Q1
Mag|$\theta$ 5 mas 10 mas 15 mas 30 mas 50 mas 120 mas
1 0.57 3.43 3.63 3.78 3.84 3.84
3 0.57 3.40 3.60 3.74 3.80 3.81
5 0.55 3.29 3.48 3.63 3.68 3.69
7 0.48 2.90 3.07 3.19 3.24 3.25
9 0.30 1.78 1.88 1.96 1.99 1.99
10 0.16 0.94 1.00 1.04 1.06 1.06
Table 2: End-to-end GMagAO-X planet throughput in percentages (from 0% to 100%) for different host star magnitudes and planet-star angular separations ($\theta$). This table is for a Max Speed AO control system and Q1 observing conditions. For different AO control systems and observing conditions see Table 5 in Appendix A.
GMT, Max Speed, Q1
Mag|$\theta$ 5 mas 10 mas 15 mas 30 mas 50 mas 120 mas
1 3.06E-08 6.09E-09 5.85E-09 4.53E-09 3.51E-09 2.35E-09
3 9.79E-08 1.97E-08 1.93E-08 1.61E-08 1.32E-08 9.50E-09
5 4.01E-07 8.10E-08 8.00E-08 6.88E-08 5.83E-08 4.39E-08
7 2.26E-06 4.57E-07 4.54E-07 4.00E-07 3.47E-07 2.71E-07
9 2.12E-05 4.30E-06 4.30E-06 3.87E-06 3.43E-06 2.72E-06
10 8.87E-05 1.80E-05 1.81E-05 1.65E-05 1.49E-05 1.15E-05
Table 3: GMT achievable $F_{p}/F_{\star}$ contrast at S/N=5 for 1 hour integration times using high contrast imaging with high-resolution spectroscopy for different host star magnitudes and planet-star angular separations ($\theta$). This table is for a Max Speed AO control system and Q1 observing conditions. For different AO control systems and observing conditions see Table 5 in Appendix A.
ELT, Max Speed, 128 actuators, Q1
Mag|$\theta$ 5 mas 10 mas 15 mas 30 mas 50 mas 120 mas
1 2.55E-09 2.08E-09 1.99E-09 1.55E-09 1.25E-09 9.36E-10
3 7.12E-09 5.84E-09 5.64E-09 4.64E-09 3.90E-09 2.95E-09
5 2.48E-08 2.04E-08 1.99E-08 1.68E-08 1.43E-08 1.09E-08
7 1.03E-07 8.49E-08 8.30E-08 7.14E-08 6.17E-08 4.78E-08
9 5.72E-07 4.72E-07 4.64E-07 4.06E-07 3.57E-07 2.82E-07
10 1.62E-06 1.34E-06 1.32E-06 1.17E-06 1.03E-06 8.24E-07
Table 4: ELT achievable $F_{p}/F_{\star}$ contrast at S/N=5 for 1 hour integration times using high contrast imaging with high-resolution spectroscopy for different host star magnitudes and planet-star angular separations ($\theta$). This table is for a Max Speed AO control system, 128 deformable mirror actuators, and Q1 observing conditions. For different AO control systems, actuators, and observing conditions see Table 7 in Appendix A.
3 Survey Simulation
In order to realistically simulate the execution of an observing survey of nearby EECs, we followed the survey simulation methodology of Hardegree-Ullman et al. (2023). First, we generated 1000 simulated local neighborhood ($d<20$ pc) exoplanet populations in Bioverse based on exoplanet demographics of Bergsten et al. (2022) and computed the time it will take to probe Earth-like O2 levels for each simulated EEC. We then sorted each population of simulated exoplanets by the number of years it would take to probe O2 and calculated the median value of the first through nth planets. The results are shown in Figure 4 for optimized instrument and observing modes, which are Max Speed, Q1 for the GMT, and Max Speed, Q1, 128 actuators for the ELT. An assessment of different instrument and observing modes is given in Appendix A.
Figure 4 shows that the ELT would be able to survey $\sim$3$\times$ more planets than the GMT for a survey of similar length, approximately commensurate with the $\sim$3$\times$ more collecting area. For example, in a 10-year survey from start to finish, if all EECs have Earth-like levels of O2, the GMT could probe up to $\sim$7 planets, and the ELT could probe up to $\sim$19. Note, that this simulation did not consider that these planets would likely have overlapping observing windows, so the number of planets surveyed for O2 is an upper limit. Further observing optimization (e.g., observing exoplanets near maximum phase, or combining measurements from multiple telescopes) should be possible, but was not explored in this study. It is also important to highlight that over 85% of the simulated planets which could be probed within a 10-year survey across all survey simulations orbit M dwarfs, and about 75% of these simulated planets are within 13 pc of the Sun (Figure 5).
Figure 4: Simulation results for a survey of Earth-like concentrations of O2 at S/N=5 on exo-Earth candidates within 20 pc with the ELT (blue dots) and the GMT (orange stars). A survey of the same duration on both telescopes would allow the ELT to probe about three times more planets for O2 than the GMT. The large error bars are driven by the wide range of possible orbital parameters for our simulated exoplanets.
Figure 5: Distributions of host star $T_{\mathrm{eff}}$ (top) and distance (bottom) across all EEC survey simulations for EECs that could be probed for Earth levels of O2 within a 10-year survey. 85% of the detectable simulated planets orbit M dwarfs, with a peak near spectral type M3. About 3/4 of the detectable simulated planets are within 13 pc of the Sun. The dashed-dotted blue lines show the distribution of all bright ($G_{RP}<10$) stars within 20 pc, which highlights that the GMT and ELT will be insensitive to O2 in the atmospheres of EECs orbiting stars above $\sim$5000 K, and is most sensitive to stars within about 13 pc.
3.1 Super-Earths
Next, we carried out a simulation to explore the observability of O2 in all super-Earth exoplanets. The previous survey simulation focused on planets nearly the same size as Earth ($0.8S^{0.25}<R_{p}<1.4~{}R_{\oplus}$). While not all exoplanets smaller than $1.4~{}R_{\oplus}$ will necessarily be Earth-like, a larger population including all super-Earth exoplanets would allow us to observationally constrain the limits of Earth-like planet atmospheres. In the following simulation, we assessed the ability to probe Earth-like O2 levels on Earth to super-Earth sized planets ($0.8S^{0.25}<R_{p}<1.8~{}R_{\oplus}$), where the upper limit is guided by the exoplanet radius valley (Fulton et al., 2017; Hardegree-Ullman et al., 2020). Results from the super-Earth survey simulation are given in Figure 6. We found that about 1.5$\times$ more planets could be surveyed if the survey is broadened from Earth-sized to all Earth and super-Earth-sized planets.
Figure 6: Simulation results for a survey of Earth-like concentrations of O2 at S/N=5 on habitable zone Earth and super-Earth-sized candidates ($0.8S^{0.25}<R_{p}<1.8~{}R_{\oplus}$) within 20 pc with the ELT (blue dots) and the GMT (orange stars). Similar to the exo-Earth simulation in Figure 4, the ELT could probe about three times more planets than the GMT. Overall, a super-Earth survey could probe about 1.5 times more planets than a survey of planets up to 1.4 $R_{\oplus}$.
3.2 Known Habitable Zone Exoplanets
Within 20 pc there are five confirmed habitable zone exoplanets which would be observable from either the GMT or ELT. These exoplanets orbit stars brighter than ${G_{RP}}=10$ and have minimum masses ($M\sin i$) that make them super-Earth candidates. These are Proxima Centauri b (Anglada-Escudé et al., 2016), Ross 128 b (Bonfils et al., 2018), GJ 273 b (Astudillo-Defru et al., 2017), Wolf 1061 c (Wright et al., 2016), and GJ 667 C c (Anglada-Escudé et al., 2012). LTT 1445 A d is another nearby candidate super-Earth habitable zone exoplanet with a relatively bright host star (Lavie et al., 2023). It is worth noting that all these planets were discovered orbiting M dwarfs using the radial velocity technique, and it is possible that their mass is significantly higher than the minimum mass provided by the RV method (see e.g., Bixel & Apai, 2017). We ran a survey simulation on these planets, assuming uniform priors on all unknown or loosely constrained parameters such as albedo, inclination angle, argument of periastron, and longitude of the ascending node. Results from this simulation are shown in Figure 7. If Earth-like levels of O2 exist on these planets, it could be probed within days to a few weeks with the ELT, and within a couple of months with the GMT. For LTT 1445 A d, it will take several months, and can only be completed on the ELT within a 10 year survey. Given the fact that the existence of these planets is already established, they will likely be high-priority targets for initial exoplanet surveys with the GMT and ELT.
Figure 7: Simulation results for a survey of Earth-like concentrations of O2 at S/N=5 for five known, and one candidate (LTT 1445 A d) habitable zone exoplanets within 20 pc with the ELT (blue) and the GMT (orange). The ELT and GMT should both be able to probe for O2 on Proxima Centauri b in less than a month. LTT 1445 A d is the only planet in the simulation for which Earth-like levels of O2 could not be probed with the GMT, but it could be probed with the ELT. More precise planet and orbital parameters would minimize the uncertainties on these observing simulations. The minimum mass ($M_{\oplus}\sin i$) of each planet is given above each relatively-sized point (Faria et al., 2022; Bonfils et al., 2018; Astudillo-Defru et al., 2017; Lavie et al., 2023; Anglada-Escudé et al., 2013).
4 Hypothesis Testing
Since our survey simulations revealed the possibility of detecting O2 on several exoplanets in a 10-year survey, particularly with the ELT, in order to assess the scientific value of such surveys, we now turn to testing related population-level hypotheses. We refer the reader to Section 5 of Bixel & Apai (2021) for a full description of Bayesian hypothesis testing with Bioverse. In general, we want to assess the statistical power of a survey to test a hypothesis by running survey simulations several times in a Monte Carlo fashion under the same assumptions (but with randomly drawn parameters) and determine the fraction of simulations that validate or reject the hypothesis. Here, we propose and aim to test the “habitable zone oxygen hypothesis”: We hypothesize that habitable zone Earth-sized exoplanets will be more likely to have Earth-like levels of O2 than planets outside the habitable zone.
We must first define the statistical effect to inject into our synthetic planet population. In our model, a planet can have or not have Earth-like levels of O2. It can further be an EEC if its orbit fulfills $a_{\mathrm{inner}}<a<a_{\mathrm{outer}}$ and its radius fulfills $0.8S^{0.25}<R_{p}<1.4~{}R_{\oplus}$. The planet is considered a non-EEC otherwise, but its atmosphere may still contain Earth-like levels of oxygen. This population of oxygen-bearing non-EECs can be considered false positives, and they present a source of noise in the hypothesis testing.
We can now describe the likelihood of a planet having O2 by
$f^{\mathrm{O}_{\mathrm{2}}}=\begin{cases}f^{\mathrm{O}_{\mathrm{2}}}_{\mathrm{EEC}}&\mathrm{if\ }a_{\mathrm{inner}}<a<a_{\mathrm{outer}}\\ &\mathrm{and\ }0.8S^{0.25}<R_{p}<1.4~{}R_{\oplus}\\ f^{\mathrm{O}_{\mathrm{2}}}_{\text{non-EEC}}&\mathrm{if\ }a<a_{\mathrm{inner}}\mathrm{\ or\ }a>a_{\mathrm{outer}}\\ &\mathrm{and\ }R_{p}>0.8S^{0.25}\\ 0&\mathrm{if\ }R_{p}<0.8S^{0.25},\end{cases}$ (8)
where $f^{\mathrm{O}_{\mathrm{2}}}_{\mathrm{EEC}}$ is the fraction of EECs with Earth-like O2 levels, and $f^{\mathrm{O}_{\mathrm{2}}}_{\text{non-EEC}}$ is the fraction of non-habitable zone exoplanets with O2.
Next, we adopt the functional form of the habitable zone oxygen hypothesis, $H_{\mathrm{HZ}}^{\mathrm{O}_{\mathrm{2}}}$, as
$H_{\mathrm{HZ}}^{\mathrm{O}_{\mathrm{2}}}(a_{\mathrm{eff}})=\begin{cases}f_{\mathrm{HZ}}^{\mathrm{O}_{\mathrm{2}}}\ \ \ \mathrm{if\ }a_{\mathrm{inner}}<a_{\mathrm{eff}}<a_{\mathrm{inner}}+\Delta a\\ f_{\mathrm{HZ}}^{\mathrm{O}_{\mathrm{2}}}(f_{\text{non-HZ}}^{\mathrm{O}_{\mathrm{2}}}/f_{\mathrm{HZ}}^{\mathrm{O}_{\mathrm{2}}})\ \ \mathrm{otherwise.}\end{cases}$ (9)
The model parameter vector $\theta$ consists of the position of the inner edge of the habitable zone $a_{\mathrm{inner}}$, the width of the habitable zone $\Delta a$, the fraction of habitable zone planets with O2 $f_{\mathrm{HZ}}^{\mathrm{O}_{\mathrm{2}}}$, and the ratio of non-habitable zone planets with O2 to habitable zone planets with O2 $f_{\text{non-HZ}}^{\mathrm{O}_{\mathrm{2}}}/f_{\mathrm{HZ}}^{\mathrm{O}_{\mathrm{2}}}$.
Figure 8: Statistical power of the habitable zone oxygen hypothesis test on the ELT for different values of $\eta_{\oplus}$ and different fractions of exo-Earth candidates (EECs, left) and fractions of exo-super Earth candidates (ESECs, right) with Earth-like levels of O2. Statistical power contours for 50% and 95% are shown in black. More planets with oxygen and/or higher values for $\eta_{\oplus}$ result in a greater ability to test this hypothesis, however, the inclusion of all super-Earths does not necessarily improve this ability. We note that these figures are super-sampled to highlight differences.
In our test case, we aim to assess the statistical power to test the above hypothesis in a 10-year survey with the ELT considering different values of $f^{\mathrm{O}_{\mathrm{2}}}_{\mathrm{EEC}}$ (ten levels between 0.1 and 1) and $\eta_{\oplus}$ (five levels between 0.075 and 0.3). For simplicity, we set $f^{\mathrm{O}_{\mathrm{2}}}_{\text{non-EEC}}=1\%$. A higher value of $f^{\mathrm{O}_{\mathrm{2}}}_{\text{non-EEC}}$ would inject more false-positive O2 signals and skew our results to require fewer detections to test our hypothesis. We intentionally keep this value low because future surveys, like the one used in this hypothesis test, will likely only probe planets within the habitable zone defined either by incident stellar flux levels, or confirmed with liquid water detections on the habitable zone planets. We imposed log-uniform priors $a_{\mathrm{inner}}=[0.1,2]$, $\Delta a=[0.01,3]$, $f_{\mathrm{HZ}}=[0.01,1]$, $f_{\text{non-HZ}}/f_{\mathrm{HZ}}=[0.01,1.0]$.
For each of these realizations, our planetary populations are still subject to intrinsic stochasticity that may impact the measured diagnostic power of the survey: planetary orbits and bulk properties are randomly drawn from distributions informed by exoplanet demographics, leading to random variations of the sample. To account for this while keeping computational costs reasonable, we ran ten randomized simulations per grid cell.
Figure 8 shows the statistical power that can be achieved for different fractions of EECs with Earth-like levels of O2 and $\eta_{\oplus}$. Our ability to test the habitable zone oxygen hypothesis is contingent on more than $\sim$1/3 of EECs having Earth-like levels of O2 if EECs are relatively common, and above $\sim$1/2 for low values of $\eta_{\oplus}$. Figure 6 suggests that a 10-year survey of exo-super Earth candidates (ESECs) could probe about 1.5$\times$ more planets for Earth-like levels of O2 than a survey of only Earth-sized planets. As such, we ran the above test with the same parameters except considering ESECs rather than EECs. Figure 8 shows that considering ESECs would not necessarily improve the ability to test the habitable zone oxygen hypothesis. One potential reason for this is that in a time-limited ground-based survey, a larger target sample would take longer to survey a significant fraction and test the hypothesis. A higher-resolution hypothesis grid with more samples per cell could test this more conclusively, but that is significantly more computationally expensive.
In the previous two hypothesis tests, we considered $f^{\mathrm{O_{2}}}_{\mathrm{EEC}}$ or $f^{\mathrm{O_{2}}}_{\mathrm{ESEC}}$ for planets up to $1.4~{}R_{\oplus}$ (EEC) and $1.8~{}R_{\oplus}$ (ESEC) for different values of $\eta_{\oplus}$. Next, instead of varying $\eta_{\oplus}$, we set $\eta_{\oplus}$ = 8.5% (consistent with the most recent estimates of $\eta_{\oplus}$ for conservative habitable zone M dwarfs from Bergsten et al., 2023) and tested the effect of maximum sampled planet size on our ability to test the habitable zone oxygen hypothesis. Figure 9 shows the results of this test, indicating the maximum sampled planet size has only a small effect on the ability to test the habitable zone oxygen hypothesis between $\sim$1.2 $R_{\oplus}$ and 1.8 $R_{\oplus}$.
Figure 9: Statistical power of the habitable zone oxygen hypothesis test on the ELT for different values of maximum sampled planet size and different fractions of planet candidates with Earth-like levels of O2. Statistical power contours for 50% and 95% are shown in black. The maximum sampled planet size has minor effect on the statistical power between $\sim$1.2 $R_{\oplus}$ and $\sim$1.8 $R_{\oplus}$. We note that this figure is super-sampled.
5 Discussion
The detection of O2 in the atmospheres of Earth-analogs, if found, will no doubt have profound impacts on astronomy and astrobiology. Our simulations show that Earth-like O2 levels could be probed on EECs with the GMT and ELT with planned instrumentation. There are already a half-dozen known probable super-Earths which may be ideal candidates to start searching for oxygen (Section 3.2), and planets like Proxima Centauri b could be high-priority science targets for the GMT and ELT. Below we discuss simulation limitations and survey feasibility, O2  other biosignatures, and the importance of population-level biosignature interpretation, and end with a brief discussion of the TMT.
5.1 Limitations and Survey Feasibility
Our simulations are based on loosely constrained (e.g., $\eta_{\oplus}$) or unknown astrophysical parameters (e.g., orbital parameters). These contribute significantly to the uncertainties of our survey simulations. Exoplanet occurrence rate studies with TESS, Roman, and PLATO, and radial velocity surveys to find new EECs around nearby stars could help further reduce survey uncertainties and find new candidates for biosignature searches.
It is important to note that the GMT and ELT will be mostly sensitive to exoplanets orbiting M dwarf stars (Figure 5), and particularly those in the habitable zone. These exoplanets will not be accessible to the Habitable Worlds Observatory, which is expected to primarily focus on Sun-like FGK stars (Mamajek & Stapelfeldt, 2024). The nature of the habitability of M dwarf exoplanets is extensively studied in the literature (e.g., Shields et al., 2013, 2016; Meadows et al., 2018a; Airapetian et al., 2020; Lobo et al., 2023) and is beyond the scope of this paper.
Since the extreme-AO systems for the ELTs are still in their planning stages, we rely on instrument parameters that are not yet finalized or that are based on current systems. Consequentially, our simulations can also be used to provide guidance to instrument teams as to the requirements for biosignature searches with the ELTs.
In Section 3, we did not consider that multiple target planets could be observable at the same time, or by multiple ELTs. We also used the median calculated contrast ratio for each simulated planet (using only contrast ratios for times the planets were observable in one year of simulated observations). For simulation purposes, these are reasonable assumptions to provide upper limits to yields for an exoplanet biosignature survey for individual telescopes.
Our 10-year survey assumed we would observe during Q1 observing conditions. Since the ELTs are not solely dedicated to exoplanet science, it is unlikely that all Q1 observing time would be dedicated to the search for biosignatures. Though, exoplanets are not always observable or at ideal phases for direct imaging during Q1 observing conditions. As we assess in Appendix A, it makes little difference whether a survey is conducted in Q1 or Q2 observing conditions, which would likely be a factor in designing a biosignature survey. Until additional nearby EECs are identified, we can only report upper limits to survey yields for a hypothetical observing scenario as presented here.
5.2 O2 as an Atmospheric Biosignature
Life on Earth has a complex evolutionary history with O2. About 2.4 billion years ago, cyanobacteria took in solar energy and produced O2 as a waste product of oxygenic photosynthesis. Eventually, enough oxygen had accumulated in the ocean to begin escaping into the atmosphere during the “Great Oxidation Event,” ushering in the Proterozoic eon and the evolution of complex life. A second great atmospheric accumulation of oxygen occurred between about 800 and 600 million years ago, leading to present-day levels of oxygen and animal life (Canfield, 2005; Kump, 2008; Sessions et al., 2009; Lyons et al., 2014, 2021). In this study, we focused on the ability to probe present-day Earth-like levels of O2 on rocky exoplanets. From 600 Mya to present day, Earth oxygen levels fluctuated between $\sim$10 and 30%, settling at $\sim$21today (Costa et al., 2014). Chaloner (1989) suggest that O2 concentrations of $\sim$35are the upper limit to sustain terrestrial vegetation due to the increased potential for combustibility. We leave the exploration of the limits of O2 detectability to future studies.
5.3 Other Atmospheric Biosignatures
O2 on Earth is almost exclusively produced by oxygenic photosynthesis (Segura et al., 2005), and its present-day abundance, which is difficult to replicate by individual abiotic mechanisms (Meadows, 2017), makes it a favorable biosignature to search for in the atmospheres of exoplanets. However, there are still mechanisms that could produce an oxygen-rich atmosphere, such as water photolysis on a pure H2O world (Wordsworth & Pierrehumbert, 2014). This necessitates robust elimination of false-positive abiotic O2 production scenarios. One way to do this is by observing multiple biosignatures (Meadows et al., 2018b).
CH4 is another exoplanet biosignature candidate, as its primary production mechanism on Earth is from methanogenic bacteria in anaeorbic environments (Segura et al., 2005; Thauer & Shima, 2008). Abiotic CH4 can be produced by volcanic and geothermal processes, or by gas-water-rock reactions (Etiope & Sherwood Lollar, 2013; Guzmán-Marmolejo et al., 2013), but on Earth, biological production of CH4 outweighs abiotic production by an order of magnitude (Kelley et al., 2005; Segura et al., 2005). The thermodynamic disequilibrium of CH4 and O2 on Earth (Hitchcock & Lovelock, 1967; Sagan et al., 1993), makes the disequilibrium pair a particularly enticing biosignature. Additionally, for M dwarf planets, which will be the primary targets for biosignature searches via direct imaging with the GMT and ELT, the photochemical lifetime of CH4 is $\sim$200 years, compared to $\sim$10 years on Earth (Segura et al., 2005; Arney, 2019), which might make it easier to detect on such planets.
CO2 and CH4 are another disequilibrium pair that may indicate life. This pair is particularly important for exoplanets with an Archean-like atmospheric composition when those two species dominated in the O2 deficient environment (Haqq-Misra et al., 2008). CH4 could be produced abiotically via mantle outgassing, but this would also produce abundant CO (Krissansen-Totton et al., 2018). A biologically active planet would likely have depleted CO levels because this would be the preferred energy source for methanogens over CO2, making the absence of CO a biological check on the CO2–CH4 pair (Zahnle et al., 2011; Krissansen-Totton et al., 2018).
Figure 10 shows the habitable zone accessibility of the biosignatures O2, CH4, CO2, and CO at red-optical and near-infrared wavelengths for an inner-working angle of $2\lambda/D$ for the GMT, TMT, ELT, and JWST/HWO at distances of 5 and 10 pc. We have also included H2O, which is the basis of our definition of a habitable zone planet (Kopparapu et al., 2014). Our study focused on the O2 A-band at 0.76 $\mu$m, which could be probed in the atmospheres of most inner habitable zone planets orbiting stars up to M4 V by the GMT, TMT, and ELT out to $\sim$10 pc. Toward near-infrared wavelengths, the inner working angle decreases, diminishing the ability to probe some bands of biosignatures such as CO and CO2 in the atmospheres of exoplanets orbiting later-type and more distant stars. We leave the exploration of the ability to probe other biosignatures with the GMT, TMT, and ELT to future studies.
Figure 10: The inner working angles at $2\lambda/D$ for 0.6–2 $\mu$m are plotted for a $\sim$6-meter JWST or anticipated HWO diameter aperture (orange solid line), the GMT (blue solid line), the TMT (green solid line), and the ELT (yellow solid line). The wavelengths of H2O (blue dashed), O2 (green dashed dotted), CH4 (orange dotted), CO2 (pink dashed), and CO (black dotted) for this wavelength region are plotted as vertical lines. At distances of 5 pc (left) and 10 pc (right), The shaded horizontal regions show the conservative habitable zones for representative G2 V through M8 V stars. If a habitable zone region falls above the inner working angle of a telescope, an exoplanet and its biosignatures should be observable at that wavelength range. For example, below 1 $\mu$m, the GMT should be able to observe O2, CH4, and H2O for planets in the habitable zones of a hypothetical M4 V star at 5 pc. The GMT will only be able to access biosignatures between 1 and 2 $\mu$m if planets fall in certain regions of the habitable zone of this hypothetical star, and it will not be able to access biosignatures beyond 2 $\mu$m.
5.4 A Statistical Interpretation of Biosignatures
Even if it is possible to probe multiple biosignatures on an exoplanet, interpretation of detections/non-detections on a single planet will vary (see e.g., Madhusudhan et al., 2023; Wogan et al., 2024; Tsai et al., 2024). Therefore, it is crucial to survey a large number of planets to study biotic and abiotic processes and interpret biosignatures statistically (Apai et al., 2019). Here, we focused on the potential to probe Earth-like levels of O2 on a statistical level on hypothetical nearby exoplanets using direct imaging and high-resolution spectroscopy. Studying exoplanets on a population level with large-aperture ground-based telescopes will allow testing of the habitable zone oxygen hypothesis and other statistical trends connected to small-planet atmosphere diversity and planet formation processes.
5.5 Thirty Meter Telescope
The TMT is the only 25–40-meter class telescope planned for the northern hemisphere. It should be noted that there are about 9,700 stars within 20 pc of the Sun brighter than 10th magnitude above a declination of 30. This is about 25% of bright, nearby stars which are inaccessible to the GMT and ELT. A northern hemisphere large-aperture telescope is critical to maximize the search for biosignatures. Additionally, about 2/3 of the targets observable from a large-aperture telescope at 20 or 30 latitude overlap with the targets observable from the GMT and ELT in the southern hemisphere.
Having multiple facilities capable of searching for biosignatures is advantageous. When we are looking for S/N=5 biosignature detections over years of observations and hundreds of datasets, it will be important to have independent verification from, e.g., different telescopes and processing techniques to minimize false-positive scenarios. Multiple facilities can both expedite data collection on individual targets and potentially expand the target pool amenable to biosignature surveys if data from different instruments can be combined, and since weather conditions at one facility become less of an observing bottleneck. Hoeijmakers et al. (2018) successfully demonstrated combining more than 2,000 high-resolution reflected light spectra from different 2–8-meter class telescopes to measure the albedo of the hot Jupiter $\tau$ Boötis b and achieved a planet-to-star contrast of $\sim$10-5. The GMT, TMT, and ELT will be able to build upon these methods to reach smaller planets and potentially detect biosignatures.
The Planetary Systems Imager (PSI) is a planned second-generation instrument for the TMT which will have extreme-AO and high resolution spectroscopic capabilities (Jensen-Clem et al., 2021, 2022). We did not model the potential of TMT+PSI in this paper because a detailed instrument model is not yet available. However, assuming similar performance to the ELT and GMT systems, the TMT could probe a $\sim$50larger sample than the GMT and – equivalently – probe a sample $\sim$66the size of the ELT sample in a 10 year survey, commensurate with the difference in telescope collecting area.
6 Summary and Conclusions
The Bioverse survey simulation and hypothesis testing framework enables comprehensive and quantitative assessments of the capabilities of upcoming ground-based telescopes and space missions to study exoplanets. We expanded the Bioverse framework to assess the ability of the ELT and the GMT to probe Earth-like levels of O2 on nearby ($<$20 pc) habitable zone Earth-sized exoplanets via extreme-AO direct imaging and high resolution spectroscopy. The following list summarizes the main conclusions of this work:
  • Accounting for recent planet occurrence rate calculations, relative system velocities, and realistic target observability, we conducted a survey simulation to determine how many exo-Earths ($0.8S^{0.25}<R_{p}<1.4~{}R_{\oplus}$) and super-Earths ($0.8S^{0.25}<R_{p}<1.8~{}R_{\oplus}$) could be probed for Earth-like levels of O2 at S/N=5 with the GMT and ELT. In 10 years, the GMT could survey up to $\sim$7 and $\sim$10 Earths and super-Earths, respectively, while the ELT could survey up to $\sim$19 and $\sim$31 (Figures 4 and 6).
  • Six known, habitable zone super-Earth candidates (Proxima Centauri b, Ross 128 b, GJ 273 b, Wolf 1061 c, LTT 1445 A d, and GJ 667 C c) are ideal targets to search for O2. Given favorable orbital and exoplanet atmosphere parameters, Earth-like levels of O2 could be probed in a matter of about one week to a few weeks of observing for four of these planets with the ELT, and about a month with the GMT (Figure 7).
  • A 10-year survey with the ELT could determine if habitable zone exo-Earths are more likely to have O2. The diagnostic power of the survey to test this habitable zone oxygen hypothesis is sensitive to the fraction of EECs with Earth-like levels of O2 and the value of $\eta_{\oplus}$, but not very sensitive to maximum planet size (Figures 8 and 9).
  • The construction of a northern-hemisphere large-aperture telescope like the TMT would increase the potential target star sample by 25%. Multiple facilities can provide independent verification of biosignature detections, expedite data collection on individual exoplanets, and potentially expand the exoplanet target sample on which to search for biosignatures.
The upcoming 25–40-meter class telescopes are going to be the first facilities capable of robustly detecting biosignatures with extreme-AO direct imaging and high resolution spectroscopy. This method is much more promising than transmission spectroscopy in the search for O2 (Hardegree-Ullman et al., 2023), and a handful of high-priority target planets are already available. The ability to validate biosignature detections and test the habitable zone oxygen hypothesis and other biosignature hypotheses, however, depends on probing a larger sample of planets than has been currently discovered.
7 Acknowledgments
We acknowledge the work of Bryan Wang to improve the Bioverse code. KKH-U acknowledges Megan Weiner Mansfield for general discussions about M dwarf exoplanet atmospheres and Suri Rukdee for discussions about the LTT 1445 A system. We thank the anonymous referee for their helpful comments which significantly improved the clarity of this manuscript.
This material is based upon work supported by the National Aeronautics and Space Administration under Agreement No. 80NSSC21K0593 for the program “Alien Earths.” The results reported herein benefited from collaborations and/or information exchange within NASA’s Nexus for Exoplanet System Science (NExSS) research coordination network sponsored by NASA’s Science Mission Directorate.
This material is based upon High Performance Computing (HPC) resources supported by the University of Arizona TRIF, UITS, and Research, Innovation, and Impact (RII) and maintained by the UArizona Research Technologies department.
This research has made use of NASA’s Astrophysics Data System.
Author contributions: KKH-U and DA developed the project and planned its implementation. DA funded the project through the “Alien Earths” program and provided regular feedback and guidance as the project progressed. KKH-U updated the Bioverse code, performed the calculations and analysis, and drafted the manuscript. SYH created the GMT and ELT instrument models and computed contrast tables for O2. MS assisted with Bioverse code issues and hypothesis test implementation. MK provided early models of PCS performance. MK and JK gave project input regarding the ELT and hypothesis testing. KW provided guidance on practical direct imaging observations. All authors contributed suggestions and edits to the manuscript.
Appendix A Different Instrument Modes
For both the GMT and ELT, we first assessed different instrument setups and observing conditions. The GMagAO-X on GMT setup assumes a baseline number of 188 deformable mirror actuators, and we looked at two different AO control system optimization strategies: Max Speed, and Strehl Optimization. For PCS on the ELT, additional calculations were made for both a 128 or 200 actuator setup. Table 5 shows results for the different AO control systems and observing conditions for the GMT, and Table 7 shows results for different AO control systems, actuator counts, and observing conditions for the ELT. For each of these setups, we also give the end-to-end planet throughput for different host star magnitudes and angular separations for the GMT/GMagAO-X in Table 6 and the ELT/PCS in Table 8.
[…]
Max Speed, Q1 Max Speed, Q2
Mag|$\theta$ 5 mas 10 mas 15 mas 30 mas 50 mas 120 mas 5 mas 10 mas 15 mas 30 mas 50 mas 120 mas
1 3.06E-08 6.09E-09 5.85E-09 4.53E-09 3.51E-09 2.35E-09 6.11E-08 1.27E-08 1.24E-08 9.52E-09 6.91E-09 4.04E-09
3 9.79E-08 1.97E-08 1.93E-08 1.61E-08 1.32E-08 9.50E-09 1.68E-07 3.48E-08 3.42E-08 2.76E-08 2.17E-08 1.49E-08
5 4.01E-07 8.10E-08 8.00E-08 6.88E-08 5.83E-08 4.39E-08 6.00E-07 1.25E-07 1.24E-07 1.07E-07 8.90E-08 6.54E-08
7 2.26E-06 4.57E-07 4.54E-07 4.00E-07 3.47E-07 2.71E-07 3.26E-06 6.74E-07 6.76E-07 6.00E-07 5.19E-07 4.02E-07
9 2.12E-05 4.30E-06 4.30E-06 3.87E-06 3.43E-06 2.72E-06 3.39E-05 6.99E-06 7.06E-06 6.44E-06 5.75E-06 4.62E-06
10 8.87E-05 1.80E-05 1.81E-05 1.65E-05 1.49E-05 1.15E-05 1.71E-04 3.52E-05 3.57E-05 3.32E-05 3.03E-05 2.40E-05
Mag Strehl Optimization, Q1 Strehl Optimization, Q2
1 3.88E-08 7.69E-09 7.33E-09 5.43E-09 3.92E-09 2.32E-09 6.11E-08 1.27E-08 1.24E-08 9.52E-09 6.91E-09 4.04E-09
3 1.46E-07 2.90E-08 2.77E-08 2.05E-08 1.48E-08 8.67E-09 2.07E-07 4.32E-08 4.22E-08 3.25E-08 2.38E-08 1.45E-08
5 7.34E-07 1.45E-07 1.38E-07 1.01E-07 7.15E-08 3.82E-08 1.04E-06 2.17E-07 2.11E-07 1.61E-07 1.15E-07 6.25E-08
7 2.53E-06 5.01E-07 4.77E-07 3.52E-07 2.53E-07 1.46E-07 4.08E-06 8.51E-07 8.30E-07 6.33E-07 4.55E-07 2.56E-07
9 1.34E-05 2.65E-06 2.52E-06 1.86E-06 1.32E-06 7.38E-07 1.83E-05 3.81E-06 3.72E-06 2.86E-06 2.09E-06 1.26E-06
10 2.30E-05 4.55E-06 4.35E-06 3.29E-06 2.48E-06 1.65E-06 5.29E-05 1.10E-05 1.08E-05 8.23E-06 5.96E-06 3.35E-06
Table 5: GMT achievable $F_{p}/F_{\star}$ contrast at S/N=5 for 1 hour integration times using high contrast imaging with high-resolution spectroscopy for different host star magnitudes and planet-star angular separations ($\theta$). Tables are given for different AO control systems (Max Speed and Strehl Optimization) and Q1 and Q2 observing conditions.
Max Speed, Q1 Max Speed, Q2
Mag|$\theta$ 5 mas 10 mas 15 mas 30 mas 50 mas 120 mas 5 mas 10 mas 15 mas 30 mas 50 mas 120 mas
1 0.57 3.43 3.63 3.78 3.84 3.84 0.56 3.35 3.55 3.69 3.75 3.75
3 0.57 3.40 3.60 3.74 3.80 3.81 0.55 3.30 3.50 3.64 3.69 3.70
5 0.55 3.29 3.48 3.63 3.68 3.69 0.52 3.14 3.32 3.46 3.51 3.52
7 0.48 2.90 3.07 3.19 3.24 3.25 0.43 2.59 2.74 2.85 2.89 2.90
9 0.30 1.78 1.88 1.96 1.99 1.99 0.21 1.24 1.31 1.36 1.38 1.39
10 0.16 0.94 1.00 1.04 1.06 1.06 0.08 0.48 0.51 0.53 0.53 0.54
Mag Strehl Optimization, Q1 Strehl Optimization, Q2
1 0.57 3.43 3.63 3.78 3.84 3.84 0.56 3.35 3.55 3.69 3.75 3.75
3 0.57 3.41 3.61 3.76 3.82 3.82 0.55 3.31 3.50 3.65 3.70 3.71
5 0.56 3.36 3.55 3.70 3.76 3.76 0.54 3.22 3.41 3.55 3.60 3.61
7 0.54 3.23 3.42 3.56 3.62 3.62 0.50 2.99 3.17 3.29 3.35 3.35
9 0.49 2.92 3.09 3.22 3.27 3.27 0.41 2.47 2.62 2.73 2.77 2.77
10 0.44 2.64 2.80 2.91 2.96 2.96 0.34 2.06 2.18 2.26 2.30 2.30
Table 6: End-to-end GMagAO-X planet throughput in percentages (from 0% to 100%) for different host star magnitudes and planet-star angular separations ($\theta$). Tables are given for different AO control systems (Max Speed and Strehl Optimization) and Q1 and Q2 observing conditions.
Max Speed, 128 actuators, Q1 Max Speed, 128 actuators, Q2
Mag|$\theta$ 5 mas 10 mas 15 mas 30 mas 50 mas 120 mas 5 mas 10 mas 15 mas 30 mas 50 mas 120 mas
1 2.55E-09 2.08E-09 1.99E-09 1.55E-09 1.25E-09 9.36E-10 5.00E-09 4.07E-09 3.89E-09 3.11E-09 2.58E-09 1.97E-09
3 7.12E-09 5.84E-09 5.64E-09 4.64E-09 3.90E-09 2.95E-09 1.30E-08 1.06E-08 1.02E-08 8.33E-09 7.06E-09 5.50E-09
5 2.48E-08 2.04E-08 1.99E-08 1.68E-08 1.43E-08 1.09E-08 3.91E-08 3.21E-08 3.12E-08 2.64E-08 2.26E-08 1.75E-08
7 1.03E-07 8.49E-08 8.30E-08 7.14E-08 6.17E-08 4.78E-08 1.49E-07 1.22E-07 1.19E-07 1.02E-07 8.82E-08 6.85E-08
9 5.72E-07 4.72E-07 4.64E-07 4.06E-07 3.57E-07 2.82E-07 7.95E-07 6.54E-07 6.41E-07 5.60E-07 4.91E-07 3.90E-07
10 1.62E-06 1.34E-06 1.32E-06 1.17E-06 1.03E-06 8.24E-07 2.34E-06 1.93E-06 1.89E-06 1.67E-06 1.49E-06 1.20E-06
Mag Max Speed, 200 actuators, Q1 Max Speed, 200 actuators, Q2
1 2.24E-09 1.83E-09 1.72E-09 1.27E-09 9.48E-10 6.18E-10 4.06E-09 3.33E-09 3.13E-09 2.27E-09 1.66E-09 1.01E-09
3 6.39E-09 5.24E-09 5.03E-09 4.02E-09 3.27E-09 2.31E-09 1.07E-08 8.77E-09 8.32E-09 6.31E-09 4.92E-09 3.39E-09
5 2.34E-08 1.93E-08 1.87E-08 1.56E-08 1.31E-08 9.73E-09 3.39E-08 2.79E-08 2.70E-08 2.21E-08 1.83E-08 1.33E-08
7 1.07E-07 8.78E-08 8.58E-08 7.37E-08 6.35E-08 4.89E-08 1.46E-07 1.20E-07 1.17E-07 9.95E-08 8.50E-08 6.46E-08
9 7.13E-07 5.88E-07 5.77E-07 5.07E-07 4.46E-07 3.52E-07 9.95E-07 8.22E-07 8.07E-07 7.07E-07 6.21E-07 4.93E-07
10 2.22E-06 1.83E-06 1.80E-06 1.60E-06 1.42E-06 1.13E-06 3.42E-06 2.82E-06 2.78E-06 2.48E-06 2.21E-06 1.78E-06
Mag Strehl Optimization, 128 actuators, Q1 Strehl Optimization, 128 actuators, Q2
1 2.55E-09 2.08E-09 1.99E-09 1.55E-09 1.25E-09 9.36E-10 5.00E-09 4.07E-09 3.89E-09 3.11E-09 2.58E-09 1.97E-09
3 8.29E-09 6.78E-09 6.44E-09 4.94E-09 3.91E-09 2.86E-09 1.57E-08 1.28E-08 1.21E-08 9.37E-09 7.48E-09 5.46E-09
5 3.03E-08 2.48E-08 2.35E-08 1.76E-08 1.36E-08 9.60E-09 5.62E-08 4.57E-08 4.30E-08 3.20E-08 2.44E-08 1.65E-08
7 1.46E-07 1.20E-07 1.13E-07 8.14E-08 5.90E-08 3.59E-08 1.88E-07 1.53E-07 1.44E-07 1.07E-07 8.11E-08 5.57E-08
9 5.94E-07 4.85E-07 4.57E-07 3.30E-07 2.40E-07 1.49E-07 9.28E-07 7.53E-07 7.06E-07 5.08E-07 3.68E-07 2.24E-07
10 1.20E-06 9.77E-07 9.20E-07 6.70E-07 4.93E-07 3.18E-07 1.81E-06 1.47E-06 1.38E-06 9.99E-07 7.34E-07 4.70E-07
Mag Strehl Optimization, 200 actuators, Q1 Strehl Optimization, 200 actuators, Q2
1 2.91E-09 2.37E-09 2.23E-09 1.59E-09 1.13E-09 6.42E-10 4.06E-09 3.33E-09 3.13E-09 2.27E-09 1.66E-09 1.01E-09
3 1.09E-08 8.92E-09 8.37E-09 5.96E-09 4.21E-09 2.33E-09 1.36E-08 1.11E-08 1.05E-08 7.58E-09 5.50E-09 3.35E-09
5 3.68E-08 3.01E-08 2.82E-08 2.02E-08 1.44E-08 8.38E-09 5.10E-08 4.18E-08 3.93E-08 2.83E-08 2.04E-08 1.21E-08
7 1.40E-07 1.14E-07 1.07E-07 7.72E-08 5.56E-08 3.33E-08 2.56E-07 2.10E-07 1.97E-07 1.40E-07 9.86E-08 5.33E-08
9 7.10E-07 5.79E-07 5.44E-07 3.88E-07 2.76E-07 1.56E-07 1.06E-06 8.70E-07 8.18E-07 5.84E-07 4.13E-07 2.29E-07
10 1.53E-06 1.25E-06 1.17E-06 8.40E-07 5.99E-07 3.46E-07 2.19E-06 1.79E-06 1.69E-06 1.21E-06 8.64E-07 4.99E-07
Table 7: ELT achievable $F_{p}/F_{\star}$ contrast at S/N=5 for 1 hour integration times using high contrast imaging with high-resolution spectroscopy for different host star magnitudes and planet-star angular separations ($\theta$). Tables are given for different AO control systems (Max Speed and Strehl Optimization), 128 and 200 deformable mirror actuators, and Q1 and Q2 observing conditions.
Max Speed, 128 actuators, Q1 Max Speed, 128 actuators, Q2
Mag|$\theta$ 5 mas 10 mas 15 mas 30 mas 50 mas 120 mas 5 mas 10 mas 15 mas 30 mas 50 mas 120 mas
1 2.74 3.88 3.92 3.95 3.96 3.96 2.32 3.29 3.33 3.36 3.36 3.36
3 2.73 3.87 3.92 3.95 3.95 3.95 2.32 3.28 3.32 3.35 3.35 3.35
5 2.72 3.86 3.90 3.94 3.94 3.94 2.30 3.26 3.30 3.33 3.33 3.34
7 2.69 3.80 3.85 3.88 3.88 3.89 2.26 3.20 3.24 3.26 3.27 3.27
9 2.48 3.51 3.55 3.59 3.59 3.59 2.00 2.84 2.87 2.90 2.90 2.90
10 2.13 3.01 3.05 3.07 3.08 3.08 1.57 2.22 2.25 2.27 2.27 2.27
Mag Max Speed, 200 actuators, Q1 Max Speed, 200 actuators, Q2
1 2.94 4.16 4.21 4.25 4.25 4.25 2.74 3.88 3.93 3.96 3.97 3.97
3 2.93 4.15 4.20 4.24 4.24 4.24 2.73 3.87 3.92 3.95 3.95 3.95
5 2.91 4.12 4.17 4.21 4.21 4.21 2.70 3.83 3.88 3.91 3.91 3.91
7 2.82 4.00 4.04 4.08 4.08 4.08 2.58 3.65 3.70 3.73 3.73 3.73
9 2.37 3.36 3.40 3.43 3.43 3.43 1.92 2.73 2.76 2.78 2.79 2.79
10 1.88 2.66 2.69 2.71 2.72 2.72 1.26 1.78 1.80 1.82 1.82 1.82
Mag Strehl Optimization, 128 actuators, Q1 Strehl Optimization, 128 actuators, Q2
1 2.74 3.88 3.92 3.95 3.96 3.96 2.32 3.29 3.33 3.36 3.36 3.36
3 2.73 3.87 3.92 3.95 3.95 3.95 2.32 3.28 3.32 3.35 3.35 3.35
5 2.73 3.86 3.91 3.94 3.94 3.95 2.31 3.27 3.31 3.34 3.34 3.34
7 2.71 3.84 3.89 3.92 3.92 3.92 2.28 3.24 3.27 3.30 3.30 3.31
9 2.67 3.78 3.82 3.86 3.86 3.86 2.22 3.14 3.18 3.20 3.21 3.21
10 2.62 3.71 3.76 3.79 3.79 3.79 2.14 3.04 3.07 3.10 3.10 3.10
Mag Strehl Optimization, 200 actuators, Q1 Strehl Optimization, 200 actuators, Q2
1 2.94 4.16 4.21 4.25 4.25 4.25 2.74 3.88 3.93 3.96 3.97 3.97
3 2.93 4.15 4.20 4.24 4.24 4.24 2.73 3.87 3.92 3.95 3.96 3.96
5 2.92 4.14 4.19 4.22 4.23 4.23 2.71 3.85 3.89 3.92 3.93 3.93
7 2.89 4.09 4.14 4.18 4.18 4.18 2.66 3.77 3.81 3.85 3.85 3.85
9 2.80 3.97 4.02 4.06 4.06 4.06 2.51 3.56 3.60 3.64 3.64 3.64
10 2.71 3.84 3.88 3.92 3.92 3.92 2.35 3.33 3.37 3.40 3.40 3.40
Table 8: End-to-end ELT/PCS planet throughput in percentages (from 0% to 100%) for different host star magnitudes and planet-star angular separations ($\theta$). Tables are given for different AO control systems (Max Speed and Strehl Optimization), 128 and 200 deformable mirror actuators, and Q1 and Q2 observing conditions.
Figure 11: Simulation results for a survey of Earth-like O2 levels at S/N=5 on exo-Earth candidates within 20 pc with different AO control systems and observing conditions for the GMT (upper left), additionally accounting for different deformable mirror actuator counts on the ELT (upper right), and considering an optimized setup for both the GMT and ELT but assessing the differences between R=100,000 and R=500,000 (lower panel).
A higher resolution spectrograph produces narrower and deeper absorption and emission lines. For O2, higher resolution gives access to more relative system velocity space. We modeled an $R=500,000$ spectrograph coupled to our AO systems, for which severe line blending in the O2 A-band occurs when $|\Delta\mathrm{RV}|<10$ km s-1 and $40\ \mathrm{km\ s}^{-1}<|\Delta\mathrm{RV}|<48\ \mathrm{km\ s}^{-1}$ (López-Morales et al., 2019; Hardegree-Ullman et al., 2023). As noted by Fowler et al. (2023), post-processing efficiency does not change at $R>100,000$, so we are able to use Tables 5 and 7 for our $R=500,000$ calculations. The lower panel of Figure 11 shows the results for $R=500,000$ compared to $R=100,000$, with a $\sim$1 or 2 planet increase in planets surveyed for a survey of 10 years, suggesting a higher resolution spectrograph is not necessarily a good investment for this science case.